# STELAR-VISION: Self-Topology-Aware Efficient Learning for Aligned Reasoning in Vision

### v3

@@ Abstract
footnotetext: Preprint
@@ 3.2 Analysis: Topological Reasoning Structures
Our analysis in Table 1 reveals that $Chain$ structure reasoning remains the best overall reasoning pattern in both VLMs, regardless of model size. However, $Tree$ and $Graph$ collectively account for more than half of the winning responses. We conduct a detailed investigation of each topology’s performance across subject domains by following the 10 subject categories in the MATH-V dataset and measuring the accuracy of each reasoning topology per subject, as illustrated in Figure 3. And thus we arrive at a conclusion that different problems benefit from different topologies, and identifying the optimal ones yields accuracy gains. This aligns with theoretical expectations—while straightforward tasks are adequately addressed by CoT-style reasoning, more complex or structurally intricate problems require richer topological representations to capture their underlying relationships effectively.
Topology-Wise Generation Length We further investigate the reasoning generation token length distribution within the TopoAug dataset. As illustrated in Figure 4, while most generations exhibit an average length of approximately 550 tokens, the $Chain$ topology produces the longest generations with a right-skewed distribution that favors extended reasoning processes. In contrast, both $Tree$ and $Graph$ topologies demonstrate similar token length distributions with overall a lower length, indicating comparable reasoning verbosity across these structured approaches.
Figure 4: Distribution of generated reasoning token length of ChainChain, TreeTree, and $Graph$ topological structures in TopoAug Dataset. The box within each violin plot represents the median, and 25% and 75% percentile thresholds.
@@ 4.1 Experimental Setup
Our base model is Qwen2VL-7B-Instruct (Wang et al. 2024b), chosen for its stable generation of all three topologies in 74% of cases. We exclude the newer Qwen2.5VL-7B-Instruct (Team 2024) for its instability in producing $Tree$ and $Graph$ structure reasoning. We also compare our models with additional open-source and proprietary models.
@@ 4.4 Efficiency Gains from Frugal Learning
We assess the impact of Frugal Learning on both in-distribution (ID) and out-of-distribution (OOD) performance, as shown in Table 6. The difference between STELAR-Vision-Short and STELAR-Vision-Short is detailed in Section 3.3.

### v4

@@ Abstract
footnotetext: This paper has been accepted at AAAI 2026. This is the author’s extended version. The final version will appear in the official proceedings.
@@ 3.2 Analysis: Topological Reasoning Structures
Our analysis in Table 1 reveals that ChainChain structure reasoning remains the best overall reasoning pattern in both VLMs, regardless of model size. However, TreeTree and GraphGraph collectively account for more than half of the winning responses. We conduct a detailed investigation of each topology’s performance across subject domains by following the 10 subject categories in the MATH-V dataset and measuring the accuracy of each reasoning topology per subject, as illustrated in Figure 3. And thus we arrive at a conclusion that different problems benefit from different topologies, and identifying the optimal ones yields accuracy gains. This aligns with theoretical expectations—while straightforward tasks are adequately addressed by CoT-style reasoning, more complex or structurally intricate problems require richer topological representations to capture their underlying relationships effectively.
Topology-Wise Generation Length We further investigate the reasoning generation token length distribution within the TopoAug dataset. As illustrated in Figure 4, while most generations exhibit an average length of approximately 550 tokens, the ChainChain topology produces the longest generations with a right-skewed distribution that favors extended reasoning processes. In contrast, both TreeTree and GraphGraph topologies demonstrate similar token length distributions with overall a lower length, indicating comparable reasoning verbosity across these structured approaches.
Figure 4: Distribution of generated reasoning token length of ChainChain, TreeTree, and GraphGraph topological structures in TopoAug Dataset. The box within each violin plot represents the median, and 25% and 75% percentile thresholds.
@@ 4.1 Experimental Setup
Our base model is Qwen2VL-7B-Instruct (Wang et al. 2024b), chosen for its stable generation of all three topologies in 74% of cases. We exclude the newer Qwen2.5VL-7B-Instruct (Team 2024) for its instability in producing TreeTree and GraphGraph structure reasoning. We also compare our models with additional open-source and proprietary models.
@@ 4.4 Efficiency Gains from Frugal Learning
We assess the impact of Frugal Learning on both in-distribution (ID) and out-of-distribution (OOD) performance, as shown in Table 6. The difference between STELAR-Vision-Short and STELAR-Vision-Short is detailed in Section $3.$